OpenStudy (ksaimouli):

solve x/x-2+3x/x-4=32-2x/x^2-6x+8

OpenStudy (ksaimouli):

plz

OpenStudy (anonymous):

\[\frac{x}{x-2}+\frac{3x}{x-4}=\frac{32-2x}{x^2-6x+8}\] like that ?

jhonyy9 (jhonyy9):

x^2 -6x +8=(x-2)(x-4) x(x-4) +3x(x-2) 2(16-x) -------------- = ---------- - so restrictions will be : x not can being 2 and 4 (x-2)(x-4) (x-2)(x-4) because than these fractions are undefined x^2 -4x +3x^2 -6x 32 -2x ----------------- = --------- (x-2)(x-4) (x-2)(x-4) 4x^2 -10x 32 -2x ---------- = --------- (x-2)(x-4) (x-2)(x-4) 4x^2 -10x =32 -2x 4x^2 -8x -32=0 4(x^2 -4x -8)=0 x^2 -4x -8=0 x_1,2=(4+/- sqrt(16+32))/2 =(4+/- 4sqrt3)/2 = 2+/- 2sqrt3